[Nps] [New Logentry] RC Update
jroche at jlab.org
jroche at jlab.org
Fri Sep 1 17:35:01 EDT 2023
Logentry Text:
--
Please send questions, comments and corrections.<p>
<b>MCC update</b>
<ul>
<li>At nominal gradient (1047 MeV/pass) but not allowance. RF recovery is continuing.
<li>Beam is available to us Tuesday, 9/05 swing shift. And as of right now, we intend to take it.
</ul>
<b> Rough timeline</b>
<ul>
<li>Frid (09/01)
<ul>
<li>Afternoon: MCC Hot check out: Hall closed from 1 to 6 p.m. Hall re-opens after 6 p.m.
</ul>
<li>Mon – Labor day
<ul>
<li>We will have a RC meeting
</ul>
<li>Ties day shift (09/05)
<ul>
<li>Morning: Hall C checklist
<li>Finish installing NPS correctors
<li>Move NPS to nominal positions (NPS at 9.5 m and 17 deg)
<li>Condensate LD2 and LH2 targets.
</ul>
<li>Tues swing
<ul>
<li>Hall to controlled access
<li>MCC to demonstrate 50 uA on the dump (start physics time)
<li>Establish beam position (including sweep magnet deflection: run plan 2.1.1, 2.1.2, and 2.1.3, Dave G)
<li>Follow run plan (Charles) - HMS detector checkout (run plan 2.1.6) - trigger checkout (plan is on the work)
</ul>
<li>Wed day
<ul>
<li>Target practical training (Silviu)
<li>Possible MCC maintenance day
</ul>
</ul>
<b>Hall status update</b>
<ul>
<li> <a href="https://logbooks.jlab.org/entry/4172735">ESR Cryo underload</a>. We are trying to use as much coolant as we can.
<ul>
<li>Cryo/vacuum issues prevent the remote (and on-demand) motion of
HMS. The HMS is parked at 16.44 deg where we want it for the start
of the data taking. SHMS motion is no problem.
<li>We will decide Tu if we need 12h access Wed to reconfigure the SHMS to help with this ESR underload.
<li>To help solve the cryo overload Target cool down to Helium in
progress. At the meeting, I (JR) mistakenly stated it had been completed.
</ul>
<li>HMS checklist (Bill started this)
<li>NPS calo checklist (tuesd)
<li>NPS magnets: complete installation on Tuesday (at least all day shift)
<li>Shifter/config documentation: please continue to update/review the <a href="https://hallcweb.jlab.org/wiki/index.php/Main_Page#Fall_2023_NPS_experiments" >Hall C wiki page</a>.
<li>Calo status
<ul>
<li>new holes drilled. The calo can reach 8.5 and 9.5 m.
<li>Malek proposes new HV for cosmics and for photon DVCS based on his analysis of run 722.
<li>gain versus sweep magnet (with LED) – prelim result 8% diff –
<li>need to tape 5 more columns
<li>gain versus HV (with LED) -pending
<li>need to remove front plate – pending
<li>need to connect NDX cable + place dosimeter and humidity and temp sensor and instal Dryer lines
</ul>
<li> <a href="https://hallcweb.jlab.org/wiki/index.php/NPS_Run_Plan_(2023)">Run plan</a>: Charles is updating it to include the commissioning of
the trigger (as was discussed Wednesday
afternoon).
</ul>
---
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4173860
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/nps/attachments/20230901/fae56258/attachment.html>
More information about the Nps
mailing list